OpenStudy (anonymous):

So I have to find the value for y in... y - 1 = 4y-2/3 I dont know if i should start with the left or right side... but i think i should start with getting rid of the fraction??

OpenStudy (anonymous):

\[y - 1 = \frac{ 4y - 2 }{ 3 }\]

OpenStudy (swissgirl):

First step would be to get rid of the fraction .... How would we go abt that??? Any ideas?

OpenStudy (anonymous):

wait not the fraction first i dont think thats possible

OpenStudy (swissgirl):

Hahahaha you were right

OpenStudy (anonymous):

I was??

OpenStudy (swissgirl):

But how would you get rid of the fraction?

OpenStudy (anonymous):

you'd multiply three by both sides...?

OpenStudy (swissgirl):

yup

OpenStudy (swissgirl):

so we would get 3y-3=4y-2

OpenStudy (anonymous):

ooooh ok

OpenStudy (anonymous):

and then you get 3y = 4y - 5 ?

OpenStudy (swissgirl):

yes

OpenStudy (swissgirl):

we can go abt it that way lol :)

OpenStudy (swissgirl):

what wld ur next step be???

OpenStudy (anonymous):

and then divide 3 by both sides??

OpenStudy (swissgirl):

Now well the issue is we have a y on both sides of the equation So we need to bring all the y's together on one side of the equation and the values on the other side How wld we get rid of the 4y on the left hand side of the equation???

OpenStudy (anonymous):

subtract 4y so that would be -1y = -5 ?

OpenStudy (swissgirl):

That looks great ..... Now we need to get rid of the -1 in front of the y

OpenStudy (anonymous):

and then divide -1 by both sides so that would be y = 5 !!

OpenStudy (swissgirl):

YAYYYYY!!!!!!!!

OpenStudy (anonymous):

YES! Thankyou XD